. A whole number is added to 25 and the same number is subtracted from 25. The sum of the resulting numbers is (A) 0 (B) 25 (C) 50 (D) 75

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Let us assume the number be x.
From the question it is given that, number is added to 25 = x + 25
The same number is subtracted to from 25 = 25–x
Then, the sum of the resulting numbers is = (x + 25) + (25 - x)

= x + 25 + 25–x
= 50 + x–x
= 50 + 0
= 50
Hence, Option C is correct
Thanks